### v4

# ScaleGNN: Towards Scalable Graph Neural Networks via Adaptive High-order Neighboring Feature Fusion

Abstract
Graph Neural Networks (GNNs) have demonstrated impressive performance across diverse graph-based tasks by leveraging message passing to capture complex node relationships. However, when applied to large-scale real-world graphs, GNNs face two major challenges: First, it becomes increasingly difficult to ensure both scalability and efficiency, as the repeated aggregation of large neighborhoods leads to significant computational overhead; Second, the over-smoothing problem arises, where excessive or deep propagation makes node representations indistinguishable, severely hindering model expressiveness. To tackle these issues, we propose ScaleGNN, a novel framework that adaptively fuses multi-hop node features for both scalable and effective graph learning. First, we construct per-hop pure neighbor matrices that capture only the exclusive structural information at each hop, avoiding the redundancy of conventional aggregation. Then, an enhanced feature fusion strategy significantly balances low-order and high-order information, preserving both local detail and global correlations without incurring excessive complexity. To further reduce redundancy and over-smoothing, we introduce a Local Contribution Score (LCS)-based masking mechanism to filter out less relevant high-order neighbors, ensuring that only the most meaningful information is aggregated. In addition, learnable sparse constraints selectively integrate multi-hop valuable features, emphasizing the most informative high-order neighbors. Extensive experiments on real-world datasets demonstrate that ScaleGNN consistently outperforms state-of-the-art GNNs in both predictive accuracy and computational efficiency, highlighting its practical value for large-scale graph learning.
Index Terms:
Scalable graph neural network, over-smoothing, large-scale graphs.
I Introduction
Graph Neural Networks (GNNs) have become the cornerstone of modern graph representation learning [1, 2], driving advances in applications such as node classification [3], link prediction [4, 5, 6], traffic prediction [7], and recommender systems [8, 9, 10]. The essence of GNNs lies in their iterative message-passing schemes [11, 12], whereby nodes update their representations by aggregating information from their local neighborhoods. This capability enables GNNs to naturally encode node features and capture node correlations.
As digital infrastructures evolve, graphs with millions or even billions of nodes and edges have become commonplace, magnifying the necessity for GNNs that are not only expressive but also computationally scalable [13, 14, 15, 16, 17]. In response, recent research communities have developed a rich array of scalable and deep GNN architectures, with significant innovations focused on balancing model expressiveness, computational efficiency, and training feasibility. One line of research is sampling-based scalable GNN methods such as GraphSAGE [18] and Cluster-GCN [19], which attempt to control computational costs by selecting subsets of neighbors during training, thereby reducing memory overhead and enabling mini-batch processing. These approaches help alleviate the burden of large-scale aggregation but may lose important structural information due to incomplete sampling. Another stream is pre-computation-based scalable approaches (including SGC [20], SIGN [21], and GAMLP [22]), which decouple neighborhood aggregation from parameter learning, allowing feature propagation to be computed in advance and greatly accelerating both training and inference[13, 14]. Meanwhile, a growing body of research has focused on deep and high-order GNNs (e.g., S2GC [23], GBP [24], RpHGNN [25]). They integrate information from multi-hop or high-order neighbors, thereby enriching the receptive field of each node and improving the capture of long-range dependencies. These advances have substantially expanded the applicability of GNNs to web-scale scenarios such as large-scale heterogeneous information networks [15, 16, 17].
Despite these substantial developments, GNNs still face two persistent and intertwined challenges when deployed on large-scale and complex graphs. First, as the number of message-passing layers increases, GNNs suffer from the over-smoothing problem that node representations become overly similar and lose their discriminative power. This issue is particularly severe when incorporating high-order neighbors [26, 27, 28, 29, 30], as their contributions may become redundant or even introduce noise, leading to performance degradation. Although several techniques, such as residual connections [31], skip connections [20, 32], and decoupling propagation [33] from feature transformation, have been proposed to alleviate over-smoothing, they often introduce additional computational complexity or fail to effectively balance local and global information. Second, many traditional GNN architectures struggle with the scalability issue that models face high consumption and computational costs when applied to large-scale graphs [34, 35, 36, 37, 38, 39, 40]. The primary reason is the exponential growth in the number of high-order neighbors, which leads to excessive information aggregation and redundant computations. Some recent works address scalability issues [25] by adopting pre-computation-based techniques, mini-batch training, or sampling strategies. While these approaches improve efficiency, they often sacrifice performance due to incomplete multi-hop neighborhood information. Recently, plug-in modules such as RMask [41] have been introduced to further enhance scalable GNNs. These modules can be seamlessly integrated into frameworks like S2GC and GAMLP, and employ techniques such as noise masking or random walk-based neighbor selection to reduce redundancy and over-smoothing. While plug-in enhancements offer practical improvements, most still fundamentally inherit the aggregation schemes of their base models [26, 27, 28, 29, 30] and often lack adaptive, effective feature fusion of information from different neighborhood hops. Overall, despite advances in sampling-based, pre-computation-based, deep and high-order aggregation, and plug-in modules, scalability and over-smoothing issues still hinder scalable GNNs for large-scale and complex graphs. A unified solution resolving both over-smoothing and scalability remains essential for accurate and efficient large-scale graph learning.
To overcome these challenges, we propose a novel framework towards scalable graph neural networks via adapting high-order neighboring feature fusion named ScaleGNN, which fuses multi-hop graph features to address the over-smoothing issue. Specifically, we introduce a learnable mechanism to construct and refine multi-hop neighbor matrices, allowing the model to adjust the relative importance of different neighborhood orders through trainable weight parameters. By regulating these weights, our approach selectively emphasizes the most informative high-order neighbors while minimizing the influence of less useful ones. Additionally, we introduce a low-order enhanced feature fusion mechanism that adaptively integrates low-order and high-order features based on their relevance to the node classification task. This ensures that our ScaleGNN effectively captures both local and global information, mitigating the over-smoothing problem. Furthermore, we incorporate a selective neighbor masking strategy named Local Contribution Score (LCS) and impose LCS-based sparse constraints to prune less important high-order neighbors, which quantify the relevance of high-order neighbors to target nodes, ensuring that only the most useful neighboring nodes contribute to feature aggregation. To validate the effectiveness of ScaleGNN, we conduct extensive experiments on multiple real-world graph datasets. The results demonstrate that our ScaleGNN consistently outperforms state-of-the-art GNN models in terms of both accuracy and efficiency, providing a scalable and robust solution to scalability and over-smoothing problems in large-scale and complex graph learning.
The main contributions of our work are as follows:
  • We design a trainable mechanism to construct and refine multi-hop neighbor matrices, effectively retaining informative high-order neighbors while filtering out redundancy and noise for better high-order information integration.
  • We propose a low-order enhanced fusion strategy that adaptively weights low-order and high-order features, enabling the model to capture both local detail and global correlations while alleviating over-smoothing.
  • We introduce the selective masking strategy named Local Contribution Score (LCS) to reduce multi-hop redundancy and impose sparse constraints based on LCS to prune unimportant high-order neighbors, significantly reducing computational cost with minimal performance loss.
  • Experiments on real-world graphs show that our model consistently outperforms state-of-the-art GNNs both in accuracy and efficiency, demonstrating superior scalability.
II Related Work
II-A Graph Neural Network
Graph Neural Networks (GNNs) are specialized neural networks for graph deep learning. GCN [42], an early model, performs message passing in each layer to aggregate 1-hop neighbor information, enriching vertex semantics. Stacking $K$ GCN layers allows GCNs to integrate information from neighbors within $K$ hops. Graph Attention Network (GAT) [43] employs attention mechanisms to prioritize important neighbors during message passing. GraphSAGE [18] enhances scalability for large graphs with neighbor sampling for mini-batch training. SGC [20] simplifies multilayer GNNs by removing nonlinearities and weight matrices between layers, leading to faster processing without sacrificing accuracy. APPNP [44] introduces an improved propagation scheme inspired by personalized PageRank [45], ensuring linear complexity via PageRank approximation. S2GC [23] uses a modified Markov Diffusion Kernel for propagation, balancing global and local vertex contexts. However, these GNNs show limitations in many real-world graphs due to challenges like heterophily [46, 47, 48] and heterogeneity. In this paper, we focus on the challenge of heterogeneity, specifically that the aforementioned GNNs are tailored for homogeneous graphs and overlook the characteristics of heterogeneous graphs [49], where vertices and edges can be of different types and contribute differently. Heterogeneous Graph Neural Networks (HGNNs) have been developed to address this by accommodating the properties of heterogeneous graphs.
HGNNs can be classified into two categories: relation-wise and representation-wise styles. Relation-wise HGNNs identify neighbors based on different relations (meta-paths) and aggregate this information for vertex representations. HAN [50] uses hand-designed meta-paths and semantic attention to enhance GAT for neighbor aggregation. MAGNN [51] involves all vertices within a meta-path for better semantics. HetGNN [52] samples neighbors via random walks and aggregates same-type neighbors with Bi-LSTMs. Representation-wise HGNNs perform multiple iterations of message passing and vertex updates. R-GCN [53] extends GCN by using distinct transformation matrices for each 1-hop relation. RSHN [54] uses a coarsened line graph for global edge-type embeddings. HetSANN [55] extends GAT with type-specific attention values. HGT [56] applies heterogeneous mutual attention with type-specific parameters. Simple-HGN [57] extends GAT by incorporating vertex features and learnable edge-type embeddings. HINormer [58] utilizes a Graph Transformer with a broader aggregation mechanism. HGAMLP [59] enhances heterogeneous graph learning by adaptively extracting and fusing local and global knowledge, while removing redundancy from high-order meta-paths, achieving state-of-the-art accuracy and efficiency on large-scale heterogeneous graphs. These HGNNs are end-to-end models but can be impractical for large graphs due to resource-intensive message passing during training.
Figure 1: The overview of the proposed ScaleGNN.
II-B Deep Graph Neural Networks
To address the scalability challenges of GNNs on large-scale graphs, deep GNNs [31, 33, 60, 61, 62] have become the dominant paradigm for graph representation learning [32, 63, 29] in recent years, owing to their powerful capability in modeling complex relational structures across various domains, such as social networks, knowledge graphs, and molecular analysis. The core idea behind GNNs is to iteratively aggregate and transform features from a node’s local neighborhood, enabling each node to capture both local and global structural information through multiple layers of message passing.
Deep GNN architectures can be broadly categorized into spectral-based and spatial-based methods. Spectral methods are grounded in graph signal processing theory and perform convolution operations in the graph spectral domain. Notably, the Graph Convolutional Network (GCN) simplifies the spectral convolution and achieves efficiency by using a localized first-order approximation, making it highly scalable for large graphs. Spatial-based methods perform aggregation directly in the graph domain, often inspired by the message passing framework. Spatial approaches have gained popularity due to their flexibility and ease of implementation for inductive learning scenarios. Additionally, Several representative deep GNN models have pushed the field forward. GraphSAGE introduces neighborhood sampling and generalizes convolution to unseen nodes, supporting inductive learning. GATs leverage attention mechanisms to adaptively weight neighbor contributions, while Graph Isomorphism Networks enhance expressive power by learning to distinguish graph structures beyond traditional message passing. In addition, more recent architectures such as PNA, GCNII, and GATv2 further improve model depth, aggregation schemes, and expressiveness.
Despite their success, deep GNNs face several significant challenges. Increasing network depth often leads to the over-smoothing problem, where node representations become indistinguishable and model performance deteriorates. Moreover, as the number of layers grows, scalability and computational complexity become major bottlenecks, especially on large-scale or dense graphs. To mitigate these issues, a variety of techniques such as residual connections, identity mapping, and attention-based neighbor selection have been proposed. In summary, deep GNN research has progressed rapidly, with continuous innovation in model architectures, aggregation mechanisms, and training paradigms. However, challenges such as over-smoothing, scalability, and effective high-order feature utilization still require further study, motivating the design of more adaptive and efficient deep GNN frameworks.
II-C Scalable Graph Neural Networks
To address the scalability challenges of GNNs on large-scale graphs, two primary research directions have emerged in scalable GNNs: sampling-based methods [64, 65, 66] and pre-computation-based techniques. While sampling reduces computation by selecting a subset of neighbors during training, it often introduces variance and information loss. In contrast, pre-computation-based approaches shift expensive operations offline, offering better stability and efficiency, and have demonstrated strong empirical performance across various benchmarks. A key idea in pre-computation is to decouple feature propagation from nonlinear transformation, allowing feature diffusion to be computed in advance without involving model parameters. SGC [20] pioneered this direction by removing nonlinearities between graph convolution layers, drastically improving scalability with minimal performance degradation. Building on this, SIGN, S2GC, GBP, and GAMLP further enhance the expressive capacity by leveraging multi-hop propagation, generalized spectral filters, or adaptive fusion, while preserving the efficiency benefits of pre-computation.
In the context of heterogeneous graphs, where multiple node and edge types introduce additional complexity, pre-computation poses new challenges. NARS [67] tackles this by sampling subsets of relations and constructing multiple subgraphs, enabling relation-specific pre-computation and reducing global information loss. SeHGNN [68] adopts a relation-wise aggregation strategy, where neighbor information is collected separately per relation within $K$ hops. This preserves finer-grained relational signals but leads to computational inefficiency as the number of relations and hops grows, particularly when $K>2$. To alleviate this trade-off, RpHGNN [69] proposes a hybrid design that combines the low information loss of the relation-wise strategy with the computational advantages of representation-wise aggregation, striking a more effective balance between accuracy and scalability in large-scale graphs.
Especially, there exist plug-in models such as RMask [41], which can be directly integrated into existing scalable GNN frameworks. These methods aim to improve both scalability and model depth by eliminating noise during feature propagation, thereby alleviating over-smoothing and reducing pre-processing overhead. By acting as flexible modules, plug-in approaches enable deeper propagation and more effective feature aggregation, significantly boosting the performance of various model-simplification GNNs across large-scale graphs.
In summary, scalable GNNs have greatly improved efficiency and applicability for large and complex graphs through sampling, pre-computation, and plug-in designs. While these methods enable faster and more scalable learning, relation-aware and adaptive aggregation further enhance flexibility for heterogeneous scenarios. However, effectively balancing scalability with information preservation, reducing sampling variance, and mitigating the over-smoothing issue remain major challenges. Continued progress relies on addressing these obstacles to achieve both scalability and high-quality representation learning.
III Preliminary
Generally, we consider an undirected large-scale graph as $\mathcal{G}=\{\mathcal{V},\mathcal{E}\,\mathcal{X}\}$, where $\mathcal{V}$ is the collection of nodes, $\mathcal{E}$ is the collection of edges between nodes, and $\mathcal{X}\in\mathbb{R}^{|\mathcal{V}|\times f}$ is the collection of node attributes, where $f$ is the size of the node feature vector. $\mathbf{A}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ denotes the adjacency matrix of graph $\mathcal{G}$. $\mathbf{D}=diag(d_{1},d_{2},\ldots,d_{|\mathcal{V}|})\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ represents the degree matrix of $\mathbf{A}$, where $d_{i}=\sum\nolimits_{{v_{j}}\in\mathcal{V}}{\mathbf{A}_{ij}}$ denotes the degree of node $v_{i}$. Given the labeled node set $\mathcal{V}_{l}$, our goal is to predict the class labels for nodes in the unlabeled set $\mathcal{V}_{u}=\mathcal{V}-\mathcal{V}_{l}$ under the supervision of $\mathcal{V}_{l}$.
IV Methodology
In this section, we present ScaleGNN, an efficient end-to-end framework for scalable and robust GNN learning. ScaleGNN addresses over-smoothing, computational inefficiency, and redundant high-order information through three synergistic components: (1) Adaptive High-order Feature Fusion, (2) Low-order Enhanced Feature Aggregation, and (3) High-order Redundant Feature Masking. These modules enable ScaleGNN to process large-scale graphs efficiently while preserving essential information across neighborhood hops. The overall framework is illustrated in Fig. 1.
IV-A Adaptive High-order Feature Fusion
Traditional GNNs are based on message passing through fixed adjacency matrices, where the adjacency structure remains constant throughout the layers. This mechanism can lead to over-smoothing, where node representations converge and lose their distinctiveness as the network depth increases. Furthermore, as nodes aggregate information from higher-order neighbors, the feature similarity across nodes can become excessive, thereby reducing the model’s expressiveness.
To mitigate these issues, we propose a method for constructing distinct adjacency matrices for each neighborhood hop. This allows us to isolate the relationships specific to each hop, ensuring that each neighborhood level contributes uniquely to the node representation. We define the pure $i$-th hop adjacency matrix $\mathbf{A}_{i}$ as the difference between the $(i-1)$-th and $i$-th hops of the adjacency matrix:
$\mathbf{A}_{i}=\mathbf{A}^{i}-\mathbf{A}^{i-1},\quad i=1,2,3,\dots,K,$ (1)
where $\mathbf{A}^{i}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ is the $i$-th hop of the adjacency matrix, which captures the connections up to $i$ hops, and $\mathbf{A}_{i}$ isolates the new relations emerging specifically at the $i$-th hop. $K$ is the maximum number of hops, also referred to as the order. Specifically, both $\mathbf{A}_{1}$ and $\mathbf{A}^{1}$ are the adjacency matrix of the graph, and $\mathbf{A}^{0}$ is the zero matrix.
Incorporating information from multiple hops can provide a more comprehensive view of the graph structure. However, the challenge lies in integrating these multiple hops in a way that preserves their distinct contributions. To address this, we introduce a learnable weight set $\boldsymbol{\alpha}=[\alpha_{1},\alpha_{2},\dots,\alpha_{K}]$, where each $\alpha_{i}\in\mathbb{R}$ ($i\in\{1,2,\ldots,K\}$) represents the specific important variant of the $i$-th hop adjacency matrix in the final aggregation:
$\tilde{\mathbf{A}}=\sum_{i=1}^{K}\alpha_{i}\mathbf{A}_{i},\quad\sum_{i=1}^{K}\alpha_{i}=1.$ (2)
These weights are learned through back-propagation, ensuring that the model dynamically adjusts the contribution of each neighborhood hop. The use of the softmax function to optimize these weights ensures that contributions across different hops remain balanced. Finally, we obtain the adaptive high-order fused embedding matrix $\mathbf{H}\in\mathbb{R}^{|\mathcal{V}|\times d}$ as follows:
$\mathbf{H}=\sigma(\tilde{\mathbf{A}}\cdot\mathcal{X}\cdot\mathbf{W}),$ (3)
where $d$ is the hidden embedding dimension, $\mathcal{X}\in\mathbb{R}^{|\mathcal{V}|\times f}$ is the input feature matrix, and $\mathbf{W}\in\mathbb{R}^{f\times d}$ is the learnable weight matrix for high-order features. The activation function $\sigma(\cdot)$ introduces non-linearity to the feature transformation.
Specifically, we use the current model as our base version named ScaleGNNb, which is relatively simple and efficient.
IV-B Low-order Enhanced Feature Aggregation
The integration of low-order (local) and high-order (global) features in many GNN-based models is done in a single and often static step. However, this can lead to high-order features dominating over low-order ones, which might diminish the model’s ability to preserve fine-grained local information. To address this, we propose explicitly separating low-order and high-order features before fusion. First, we obtain the low-order feature matrix $\mathbf{H}_{\text{low}}$ and the high-order feature matrix $\mathbf{H}_{\text{high}}$ as follows:
$\mathbf{H}_{\text{low}}=\sigma(\mathbf{A}^{2}\cdot\mathcal{X}\cdot\mathbf{W}_{\text{low}}),\quad\mathbf{H}_{\text{high}}=\sigma(\tilde{\mathbf{A}}\cdot\mathcal{X}\cdot\mathbf{W}_{\text{high}}),$ (4)
where $\mathbf{W}_{\text{low}},\mathbf{W}_{\text{high}}\in\mathbb{R}^{f\times d}$ are both learnable weight matrices for low-order and high-order features, respectively.
To ensure a balanced integration of low-order and high-order features, we introduce a learnable balancing factor $\beta$, which allows for dynamic adjustment of diverse contributions from two feature types. The final node representation $\mathbf{H}$ is the weighted sum of low-order and high-order features:
$\mathbf{H}=\beta\cdot\mathbf{H}_{\text{low}}+(1-\beta)\cdot\mathbf{H}_{\text{high}},$ (5)
where $\beta$ is optimized during training, ensuring the model can adaptively balance the trade-off between local and global information, depending on the specific task and graph structure.
This separation and fusion process enhances the model’s ability to capture both fine-grained, local details and broader, global patterns in the graph, ultimately leading to more expressive and robust node embeddings.
IV-C High-order Redundant Feature Masking
Higher-order neighbors introduce diverse, yet potentially redundant, information to GNNs. These neighbors will introduce noise without an effective masking mechanism, diluting the informative signal and impairing the model’s ability to generate useful node embeddings. To address this, we propose a novel method for selecting the most relevant and masking other redundant high-order neighbors based on their local contributions to node representations.
Inspired by the scaled dot-product attention mechanism [70], we introduce the Local Contribution Score (LCS) to assign importance scores to high-order neighbors. The LCS is designed to quantify the relevance of each neighboring node of the target node by considering both structural similarity and feature alignment. Specifically, for a node $v$ and its $i$-th hop neighbor $u$, the LCS value is computed as follows:
$\text{LCS}(v,u,i)=\frac{\exp\left(\frac{(\mathbf{W}_{1}\mathbf{x}_{v})^{\top}(\mathbf{W}_{2}\mathbf{x}_{u})}{\sqrt{d_{f}}}\right)}{\sum\limits_{u^{\prime}\in\mathcal{N}_{i}(v)}\exp\left(\frac{(\mathbf{W}_{1}\mathbf{x}_{v})^{\top}(\mathbf{W}_{2}\mathbf{x}_{u^{\prime}})}{\sqrt{d_{f}}}\right)}.$ (6)
Here, $\mathbf{x}_{v},\mathbf{x}_{u}\in\mathbb{R}^{f}$ are the input feature vectors of the target node $v$ and its $k$-hop neighbor $u$. $\mathbf{W}_{1},\mathbf{W}_{2}\in\mathbb{R}^{d_{f}\times f}$ are learnable projection matrices that map features into a latent attention space of dimension $d_{f}$, which is a tunable hyperparameter, and typically $d_{f}\leq f$ to ensure efficiency. $\mathcal{N}_{i}(v)$ denotes the $i$-hop neighborhood of node $v$ induced by $\mathbf{A}_{i}$. The denominator normalizes attention scores within each local $i$-hop neighborhood using the softmax function. This equation allows the model to learn context-aware neighbor importance scores in a differentiable and robust way.
The LCS allows us to mask redundant high-order neighbors and identify the most relevant neighbors by considering both local structure and feature similarity, enabling the model to selectively pass information from the most meaningful high-order neighbors. Once the LCS is computed for each neighbor, we select the top-$m_{i}$ neighbors with the highest LCS values for each node among its $i$-hop neighbors:
$\mathbf{A}_{i}^{\text{filter}}=\text{Top}_{m_{i}}(\mathbf{A}_{i},\text{LCS}),$ (7)
[…]
IV-D Joint Optimization
To ensure the effectiveness of our proposed ScaleGNN, we design an optimization framework that consists of three key loss components: the node classification task loss, the LCS regularization loss, and the sparse constraint loss on the high-order information. These losses work together to enhance model performance, suppress noise from uninformative high-order neighbors, and maintain computational efficiency.
For node classification, we employ the standard cross-entropy loss as the primary task objective:
$\mathcal{L}_{Cla}=-\sum_{i\in\mathcal{V}}y_{i}\log\hat{y}_{i},$ (8)
where $y_{i}$ represents the one-hot vector of the ground-truth class label for node $i$, and $\hat{y}_{i}$ is the predicted probability distribution over classes generated by a softmax layer.
To further mitigate the over-smoothing issue, we impose a sparse constraint (SC) on high-order feature propagation:
$\mathcal{L}_{SC}=\sum_{i=2}^{K}\|\bar{\mathbf{A}}_{i}^{\text{filter}}\|_{1},$ (9)
where $\bar{\mathbf{A}}_{i}^{\text{filter}}$ is the filtered high-order adjacency matrix, $\bar{\mathbf{A}}_{i}^{\text{filter}}=\mathbb{I}\left[\mathbf{A}_{i}^{\text{filter}}>0\right]$, where $\mathbb{I}[\cdot]$ denotes the element-wise indicator function, and $\|\cdot\|_{1}$ denotes the $\ell_{1}$ norm. This term constrains the model to minimize the number of neighbors retained for each node at each hop.
To guide the model in selecting meaningful high-order neighbors, we also introduce an LCS-based regularization:
$\mathcal{L}_{\text{LCS}}=\sum_{v\in\mathcal{V}}\sum_{i=2}^{K}\sum_{u\in\mathcal{N}_{i}^{\text{select}}(v)}\left(1-\text{LCS}(v,u,i)\right)^{2},$ (10)
where $\mathcal{N}_{i}^{\text{select}}(v)$ is the retained neighboring nodes which are not been filtered. This term enforces high-order neighbors to be selected based on their LCS values, reducing the influence of noisy or redundant neighbors.
Combining all components, the final optimization objective is as follows:
$\mathcal{L}=\mathcal{L}_{Cla}+\lambda_{1}\mathcal{L}_{\text{LCS}}+\lambda_{2}\mathcal{L}_{SC},$ (11)
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters balancing regularization against the primary task objective. Specifically, enforcing sparsity constraints during the selection of higher-order neighbors per hop allows the model to strike a balance between preserving critical features and mitigating redundant information.
Overall, this optimization framework ensures that our ScaleGNN maintains a balance between node classification performance, robust high-order neighbor aggregation, and computational efficiency.
Algorithm 1 ScaleGNN
0:  Graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{X})$, adjacency $\mathbf{A}$, max order $K$
0:  Node representations $\mathbf{H}$
1:  Construct pure $i$-hop adjacency:
2:  for $i=1$ to $K$ do
3:     $\mathbf{A}_{i}=\mathbf{A}^{i}-\mathbf{A}^{i-1}$
4:  end for
5:  Adaptive high-order feature fusion:
6:  Learn weights $\alpha_{1},\dots,\alpha_{K}$ via softmax, $\sum_{i}\alpha_{i}=1$
7:  $\tilde{\mathbf{A}}=\sum_{i=1}^{K}\alpha_{i}\mathbf{A}_{i}$
8:  $\mathbf{H}_{\text{high}}=\sigma(\tilde{\mathbf{A}}\cdot\mathcal{X}\cdot\mathbf{W}_{\text{high}})$
9:  Low-order feature aggregation:
10:  $\mathbf{H}_{\text{low}}=\sigma(\mathbf{A}^{2}\cdot\mathcal{X}\cdot\mathbf{W}_{\text{low}})$
11:  High-order redundant feature masking:
12:  for each node $v$ do
13:     for $i=2$ to $K$ do
14:        for each neighbor $u$ in $\mathbf{A}_{i}$ do
15:           Calculate $\text{LCS}(v,u,i)$ by Eq. (6)
16:        end for
17:        Select top-$m_{i}$ neighbors by LCS, mask others in $\mathbf{A}_{i}$
18:     end for
19:  end for
20:  Update $\tilde{\mathbf{A}}$ and $\mathbf{H}_{\text{high}}$ using $\bar{\mathbf{A}}_{i}^{\text{filter}}$
21:  Feature fusion:
22:  Learn $\beta\in[0,1]$
23:  $\mathbf{H}=\beta\mathbf{H}_{\text{low}}+(1-\beta)\mathbf{H}_{\text{high}}$
24:  return  $\mathbf{H}$
Algorithm 1 summarizes the overall process of the proposed ScaleGNN framework. The model first constructs pure $k$-hop adjacency matrices to isolate neighbor information at different topological scales. It then performs adaptive high-order feature fusion by learning attention weights across multiple hops, and separately aggregates low-order features via the original adjacency. To suppress redundant information introduced by distant neighbors, a local contribution score (LCS) is used to select the most informative nodes, refining the high-order structure. Finally, the node representations are obtained by weighted fusion of the low- and high-order embeddings.
IV-E Time Complexity Analysis
To evaluate the computational efficiency of ScaleGNN, we analyze its time complexity and compare it with existing scalable GNNs. The overall complexity consists of three main components: pre-processing, training, and inference.
Let $|\mathcal{V}|$ be the number of nodes, $|\mathcal{E}|$ be the number of edges, and $f$ be the feature dimension, $K$ be the maximum number of hops, $n$ be the maximum number of top-$m_{i}$ ($i\in\{1,2,\ldots,K\}$) neighbors. $L$ is the number of layers in MLP classifiers of SIGN, GAMLP, and GBP, $S$ in NARS denotes the number of subgraphs, $P$ in SeHGNN denotes the number of divided groups, and $R$ in RpHGNN is the number of node classes. In our proposed ScaleGNN, we perform the following major computations:
  • Multi-hop Adjacency Construction: To construct order-distilled adjacency matrices $\{\mathbf{A}_{i}\}_{i=1}^{K}$ where $\mathbf{A}_{i}=\mathbf{A}^{i}-\mathbf{A}^{i-1}$, we perform $K$ sparse matrix multiplications. The cost per multiplication is $O(|\mathcal{E}|)$ in sparse setting, so the total cost is $O(K|\mathcal{E}|)$.
  • LCS-based Neighbor Filtering: For each node $v$ and $k$-hop neighbor $u$, attention-based LCS involves two linear projections ($\mathbf{W}_{1}\mathbf{x}_{v}$, $\mathbf{W}_{2}\mathbf{x}_{u}$), a dot-product, and softmax computation: (1) Projection cost: amortized $O(|\mathcal{V}|fd_{f})$ (precomputed). (2) Attention score computation: $O(K|\mathcal{V}|nd_{f})$. (3) top-$m_{k}$ filtering: $O(K|\mathcal{V}|n\log n)$ using efficient heap sort. The overall complexity is approximately $O(Kn|\mathcal{V}|f)$ and remains linear in retained neighbors and supports batch-wise parallel execution.
  • Feature Aggregation and Training: Once filtered adjacency matrices are obtained, the time complexity of aggregation step is $O(Kn|\mathcal{V}|f)$. The time complexity of training and inference is $O(|\mathcal{V}|f^{2})$
Finally the end-to-end time complexity of our ScaleGNN is approximately $O(K|\mathcal{E}|+Kn|\mathcal{V}|f+|\mathcal{V}|f^{2}$).
Table I compares the time complexity of ScaleGNN with other scalable GNN architectures.
TABLE I: Time complexity comparison of existing deep and scalable GNN models.
Method Pre-processing Training Inference
S2GC $\mathcal{O}(K|\mathcal{E}|f)$ $\mathcal{O}(|\mathcal{V}|f^{2})$ $\mathcal{O}(|\mathcal{V}|f^{2})$
SIGN $\mathcal{O}(K|\mathcal{E}|f)$ $\mathcal{O}(L|\mathcal{V}|f^{2})$ $\mathcal{O}(L|\mathcal{V}|f^{2})$
GAMLP $\mathcal{O}(K|\mathcal{E}|f)$ $\mathcal{O}(L|\mathcal{V}|f^{2})$ $\mathcal{O}(L|\mathcal{V}|f^{2})$
GBP $\mathcal{O}(K|\mathcal{E}|f+K\sqrt{|\mathcal{E}|}\log|\mathcal{V}|)$ $\mathcal{O}(L|\mathcal{V}|f^{2})$ $\mathcal{O}(L|\mathcal{V}|f^{2})$
NARS $\mathcal{O}(K|\mathcal{V}|Sf+K|\mathcal{E}|)$ $\mathcal{O}(|\mathcal{V}|f^{2})$ $\mathcal{O}(|\mathcal{V}|f^{2})$
SeHGNN $\mathcal{O}(K|\mathcal{V}|Pf+K|\mathcal{E}|)$ $\mathcal{O}(|\mathcal{V}|f^{2})$ $\mathcal{O}(|\mathcal{V}|f^{2})$
RpHGNN $\mathcal{O}(K|\mathcal{V}|Rf+K|\mathcal{E}|)$ $\mathcal{O}(|\mathcal{V}|f^{2})$ $\mathcal{O}(|\mathcal{V}|f^{2})$
ScaleGNN $\mathcal{O}(Kn|\mathcal{V}|f)+K|\mathcal{E}|)$ $\mathcal{O}(|\mathcal{V}|f^{2})$ $\mathcal{O}(|\mathcal{V}|f^{2})$
Compared to traditional GNNs that suffer from excessive computation due to repeated feature propagation, ScaleGNN optimizes computational efficiency in two ways: (1) Sparse high-hop neighbor selection: By leveraging LCS-based filtering, ScaleGNN significantly reduces the number of unnecessary high-hop neighbors involved in feature propagation, improving computational efficiency. (2) Lightweight aggregation: Since ScaleGNN explicitly separates low-hop and high-hop features, it avoids redundant computations, leading to lower memory and runtime costs.
Empirical results on real-world datasets demonstrate that ScaleGNN achieves faster convergence and lower training time compared to existing model-simplification methods while maintaining high predictive accuracy.
TABLE II: Statistical summaries of datasets.
Datasets #Nodes #Edges #Features #Classes #Train/Val/Test Description
Citeseer 3,327 4,732 3,703 6 120/500/1000 citation network
Cora 2,708 5,429 1,433 7 140/500/1000 citation network
Pubmed 19,717 44,338 500 3 60/500/1000 citation network
ogbn-arxiv 169,343 1,166,243 128 40 91K/30K/49K citation network
ogbn-products 2,449,029 61,859,140 100 47 196K/49K/2,204K co-purchasing network
ogbn-papers100M 111,059,956 1,615,685,872 128 172 1200k/200k/146k citation network
V Experiments
In this section, we evaluate the performance of our proposed method through extensive experiments and answer the following questions:
  • (RQ1) Can ScaleGNN outperform SOTA GNN methods regarding predictive accuracy on real-world datasets?
  • (RQ2) How does the efficiency of ScaleGNN compare to other baseline methods?
  • (RQ3) How dose ScaleGNN mitigate the over-smoothing issue?
  • (RQ4) What are the effects of different modules in ScaleGNN on performance?
  • (RQ5) How does ScaleGNN achieves the trade-off between efficiency and accuracy?
  • (RQ6) How do different hyperparameter settings affect the model performance?
V-A Experimental Settings
V-A1 Datasets
We have evaluated the effectiveness of ScaleGNN using six real-world graph datasets, including Citeseer, Cora, Pubmed, and ogbn-arxiv, ogbn-products, ogbn-papers100M. The first three datasets (Citeseer, Cora, and Pubmed) are relatively small and have been adopted by existing work [15], while the latter three (ogbn-arxiv, ogbn-products, and ogbn-papers100M) are large-scale heterogeneous graphs commonly used in scalable HGNN evaluations. The detailed description of these datasets is shown in Table II.
V-A2 Baselines
We compare our method with three categories of GNN baseline methods as follows:
Traditional GNN methods:
  • R-GCN [53] introduces relation-specific graph convolution operations to effectively model multi-relational knowledge graphs and achieves strong performance on tasks like link prediction and entity classification.
  • Cluster-GCN [19] is an efficient GCN algorithm that trains models by sampling dense subgraphs partitioned via graph clustering. It significantly reduces memory and computational overhead, enabling efficient training on million-scale graphs for the first time.
  • HetGNN [52] is a heterogeneous graph neural network that jointly learns from both heterogeneous graph structure and node content by aggregating multi-type neighbor and attribute information, achieving superior performance on tasks like link prediction and node classification.
  • HAN [50] employs a hierarchical attention mechanism, combining node-level and semantic-level attention to effectively capture the importance of both meta-path neighbors and meta-paths, enabling expressive node representations in heterogeneous graphs.
  • MAGNN [51] integrates node content, intermediate nodes along meta-paths, and multiple meta-paths through hierarchical aggregation, enabling more expressive and accurate representations in heterogeneous graphs.
  • Simple-HGN [57] is a strong and straightforward baseline for heterogeneous graph learning that, with proper input processing, outperforms previous heterogeneous GNN models across standardized benchmark datasets.
  • HINormer [58] leverages a graph transformer architecture with global-range attention and specialized encoders to effectively capture both structural and heterogeneous relational information, enabling comprehensive node representations in heterogeneous information networks.
Deep GNN methods:
  • SIGN [21] eliminates the need for neighbor sampling by using efficiently precomputed graph convolutional filters of varying sizes, enabling fast and scalable training and inference on large-scale graphs.
  • S2GC [23] uses a modified Markov diffusion kernel to efficiently aggregate information from both local and global neighborhoods, capturing multi-scale context while mitigating over-smoothing in deeper graph convolutions.
  • GBP [24] employs a localized bidirectional propagation strategy to achieve efficient and scalable graph learning, enabling sub-linear time complexity and state-of-the-art performance on massive graphs without relying on sampling.
  • GAMLP [22] adaptively integrates information from multiple neighborhood scales using attention, effectively addressing over-smoothing and enabling highly scalable and efficient graph learning on large-scale and sparse graphs.
  • GRAND+ [71] enhances scalable graph learning by introducing an efficient pre-computation algorithm for propagation and a confidence-aware consistency loss, enabling fast and accurate semi-supervised learning on large-scale graphs.
  • LazyGNN [72] captures long-distance dependencies in graphs using shallow models rather than deeper architectures, effectively avoiding neighborhood explosion and achieving efficient, scalable performance on large-scale graphs.
Scalable GNN methods:
  • NARS [67] efficiently computes neighbor-averaged features over sampled relation subgraphs, enabling scalable and accurate learning on large heterogeneous graphs without complex end-to-end GNN architectures.
  • SeHGNN [68] streamlines heterogeneous graph learning by using pre-computed mean aggregation and a single-layer structure with long meta-paths, combined with transformer-based semantic fusion, achieving high accuracy and fast training with low complexity.
  • RpHGNN [69] combines efficient pre-computation with low information loss by using random projection squashing and a relation-wise even-odd propagation scheme, enabling fast and accurate learning on heterogeneous graphs.
TABLE III: Performance comparison of all models on six datasets. Mi-F1 and Ma-F1 are short for Micro-F1 and Macro-F1. Marker * indicates the results are statistically significant against the best-performed baselines (t-test with p-value $<$ 0.01).
Method Citeseer Cora Pubmed ogbn-arxiv ogbn-products ogbn-paper100M
Mi-F1 Ma-F1 Mi-F1 Ma-F1 Mi-F1 Ma-F1 Mi-F1 Ma-F1 Mi-F1 Ma-F1 Mi-F1 Ma-F1
Trad. GNN RGCN 70.53 70.14 81.23 80.59 78.37 75.58 58.26 46.83 48.94 48.04 42.31 31.12
Cluster-GCN 71.25 70.72 82.10 80.88 78.84 76.06 78.96 75.95 71.59 63.88 64.34 55.72
HetGNN 70.55 70.13 81.20 80.57 78.40 75.66 OOM. OOM. OOM. OOM. OOM. OOM.
HAN 71.10 70.57 81.39 80.86 78.53 75.85 54.68 31.50 OOM. OOM. OOM. OOM.
MAGNN 71.22 70.69 81.57 80.98 78.55 76.01 OOM. OOM. OOM. OOM. OOM. OOM.
Simple-HGN 71.34 70.78 81.88 81.14 78.88 76.17 69.47 58.29 OOM. OOM. OOM. OOM.
HINormer 72.75 71.91 82.19 81.67 79.35 76.60 70.71 59.68 OOM. OOM. OOM. OOM.
Deep GNN SIGN 72.44 71.97 82.11 80.46 79.53 76.44 71.78 62.99 78.21 70.10 64.32 55.49
S2GC 72.78 72.13 82.45 82.09 79.56 76.67 71.88 63.23 77.09 69.54 64.63 55.90
GBP 72.61 72.08 83.52 82.42 80.56 79.13 71.42 62.41 77.27 69.49 64.66 56.07
GAMLP 72.59 72.06 82.31 80.32 79.11 76.15 71.88 63.15 80.34 71.37 64.53 55.90
GRAND+ 72.48 72.03 82.04 80.23 79.24 76.46 71.43 62.30 77.52 69.70 64.11 55.19
LazyGNN 72.91 72.21 82.45 80.87 79.92 77.03 71.63 62.72 77.60 69.90 64.25 55.38
Scal. GNN NARS 72.41 71.21 81.79 80.30 78.82 75.82 70.75 59.88 78.08 70.20 63.52 54.63
SeHGNN 72.90 72.45 83.04 81.43 79.71 77.70 71.82 62.69 79.68 70.85 64.55 55.86
RpHGNN 73.12 72.54 83.27 82.37 80.60 79.32 72.06 63.34 80.89 71.44 64.95 56.50
Ours ScaleGNNb 72.51 72.74 83.42 82.60 80.58 79.25 71.98 63.11 80.75 71.34 64.97 56.60
ScaleGNN 74.44* 73.76* 84.75* 84.21* 82.11* 80.67* 73.25* 64.20* 81.82* 72.47* 65.69* 57.42*
Notice that most of the traditional and deep GNN methods encounter out-of-memory (OOM) issues when dealing with large datasets (such as products and paper100M). As for R-GCN, while the original version presents OOM issues with large datasets, it is often used as a baseline for large-scale HGNNs. To adapt R-GCN for large datasets, we follow NARS’ experimental setting and adopt the neighbor sampling strategy used by GraphSAGE [18]. Besides, scalable GNN methods, NARS, SeHGNN, and RpHGNN, are all pre-computation-based methods that differ from other end-to-end models. Most of their model runtime is consumed in obtaining the pre-computation tensor for large-scale graphs, with only a tiny amount of time spent on training and inference.
V-A3 Implementation Details
We implement all baseline methods according to their provided code. For three small datasets, we set the epoch number to 100 and select dropout rate, weight decay, and learning rate from {0.1, 0.2, 0.3, 0.4, 0.5, 0.6}, {1e-5, 1e-4, 1e-3} and {0.001, 0.005, 0.01, 0.05, 0.1}, respectively. For three large-scale datasets, we set the epoch number and learning rate to 1000 and 0.001 without weight decay. The embedding dimension $d$ is chosen from {32, 64, 128, 256, 512}, and our method uses GCN as the message aggregator.
V-B Performance Comparison (RQ1)
We evaluate the performance of our ScaleGNN and all baselines on six real-world datasets under the node classification task. The experimental results are shown in Table III. The best results are highlighted in bold, and the second-best results are underlined. As we can see, our ScaleGNN achieves the optimal performance, significantly outperforming all baselines in both micro-F1 and macro-F1 metrics on six datasets, with specific emphasis on the improvement relative to the SOTA methods.
Traditional GNN methods work flawlessly on small-scale datasets, but most of them cannot be extended to large-scale graphs and are prone to OOM issues. For example, R-GCN, HAN, Simple-HGN, and HINormer can get experimental results on ogbn-arxiv, but cannot run on larger-scale graphs such as ogbn-products and ogbn-paper100M. For deep GNN methods, they work on learning large-scale graph representations using deeper GNNs while mitigating the over-smoothing problem on large-scale graphs. A common issue with such methods is that they typically convolve dozens or even tens of GNN layers. While capturing the caveat features to some extent, it also greatly increases the computational complexity of the model. Pre-computation-based GNNs are distinguished from end-to-end models by the fact that they consume a significant amount of time in pre-computing the tensor of large-scale graphs, so that the models do not need to iteratively learn graph representations during training and inference.
It is worth mentioning that our basic model ScaleGNNb, although retaining only the basic adaptive high-order feature fusion module, achieves a very similar performance to the SOTA method RpHGNN and runs efficiently, far outperforming all baselines. ScaleGNN gains further performance despite being less efficient than ScaleGNNb. Therefore, ScaleGNN and ScaleGNNb provide an important reference value for us to explore the trade-off between model accuracy and efficiency.
V-C Efficiency Analysis (RQ2)
Figure 2: Efficiency comparison with SOTA GNN methods, and speedup analysis compared with our ScaleGNNb.
To evaluate the efficiency of our proposed methods, ScaleGNN and ScaleGNNb, we compare their total runtime with several state-of-the-art baselines, including S2GC, GBP, GAMLP, and RpHGNN, across three benchmark datasets: ogbn-arxiv, ogbn-products, and ogbn-papers100M. The results presented in Fig. 2 clearly demonstrate the superior efficiency of ScaleGNN and ScaleGNNb over all competing methods.
V-C1 Computation Efficiency
As shown in the figure, ScaleGNNb achieves the lowest runtime across all datasets, followed closely by ScaleGNN. Specifically, on the large-scale ogbn-papers100M dataset, ScaleGNNb completes the model running in only 4,110.2 seconds, significantly outperforming S2GC (19,914.4s), GBP (35,706.1s), GAMLP (35,656.6s), and RpHGNN (10,399.8s). This trend is also consistently observed across ogbn-arxiv and ogbn-products datasets, where ScaleGNNb reduces runtime by at least 4.1× to 6.7× compared to the best-performing baseline.
The significant reduction in runtime can be attributed to the efficient design of ScaleGNN, which optimizes neighborhood aggregation and selectively samples informative neighbors, thereby reducing computational overhead while preserving model performance. Moreover, the enhanced scalability of ScaleGNNb further refines this efficiency by introducing additional optimizations in feature propagation and memory management.
V-C2 Contrastive Speedup Analysis
To better illustrate the efficiency advantage, we annotate each baseline method with its relative speedup factor compared to ScaleGNNb. It is evident that methods like GBP and GAMLP exhibit substantially higher computational costs, with runtime exceeding 6× that of ScaleGNNb in certain cases. Even RpHGNN, which demonstrates competitive efficiency, remains 2.5× to 3.1× slower than ScaleGNNb on large datasets. These results highlight the capability of ScaleGNN to effectively mitigate the scalability limitations of existing methods. By reducing redundant computations and dynamically adjusting the information flow, our approach enables substantial efficiency improvements without compromising accuracy. The consistent speedup across diverse datasets further underscores the robustness of our design, making it particularly suitable for large-scale graph learning tasks.
In summary, ScaleGNN and ScaleGNNb achieve remarkable computational efficiency, substantially reducing runtime compared to existing approaches. The combination of selective neighbor aggregation, adaptive feature diffusion, and memory-efficient propagation mechanisms enables our methods to handle large-scale graphs with minimal computational overhead. These findings underscore the practical applicability of ScaleGNN in real-world scenarios where the model efficiency is a critical concern.
V-D Mitigating Over-smoothing Issue (RQ3)
Over-smoothing is a well-known issue in deep graph neural networks, where node representations become indistinguishable as the number of propagation layers increases. This problem is particularly severe in large-scale graphs such as ogbn-papers100M, where excessive feature mixing leads to performance degradation. To evaluate the effectiveness of our approach in addressing this challenge, we compare ScaleGNN with several SOTA baselines, including S2GC, GBP, GAMLP, and RpHGNN, across different propagation depths.
(a) ogbn-products
(b) ogbn-paper100M
Figure 3: Experimental results of our model over SOTA GNN models w.r.t. the number of hops.
Fig. 3 shows that traditional GNN methods experience significant performance drops as the number of propagation layers increases. For example, S2GC, GBP, and GAMLP achieve their peak performance at moderate depths (around 8-12 layers) but degrade rapidly beyond this point. This phenomenon occurs because deeper layers cause excessive aggregation, leading to the loss of discriminative node features. In contrast, ScaleGNN exhibits a more stable performance curve. Even at greater depths (e.g., 30 layers), ScaleGNN maintains a high micro-F1 score, outperforming all baselines. This robustness demonstrates that ScaleGNN effectively preserves structural and feature information, preventing the collapse of node representations.
The key to ScaleGNN’s resilience against over-smoothing lies in its adaptive feature aggregation mechanism. Unlike conventional approaches that uniformly aggregate neighborhood information, ScaleGNN dynamically balances local and global features, ensuring that useful information is retained while mitigating excessive smoothing. Specifically, our ScaleGNN selectively incorporates higher-order neighbors with controlled influence, preventing over-mixing of features. It employs adaptive diffusion mechanisms to regulate information propagation, ensuring a balance between short-term and long-term dependencies. ScaleGNN utilizes LCS to prioritize informative neighbors, reducing noise and redundancy in feature aggregation. These strategies allow ScaleGNN to maintain a richer set of node representations, even in deep architectures, making it particularly effective for large-scale graphs.
V-E Ablation Study (RQ4)
To evaluate the effectiveness of each component in ScaleGNN, we further conduct ablation studies on different variants. Specifically, we generate three variants as follows:
  • w/o Ada removes the adaptive high-order feature fusion module and employs the deep GNN method S2GC [23] to obtain the graph representation.
  • w/o Low excludes the low-order enhanced feature aggregation and only employs adaptive high-order feature fusion.
  • w/o LCS removes the high-order redundant feature masking mechanism.
  • w/o SC removes the masking sparse constraint and keeps the same retained number of nodes in each $\mathbf{A}_{k}^{\text{filter}}$.
TABLE IV: The comparison of ScaleGNN and its variants on micro-F1 and macro-F1 (Mi-F1 and Ma-F1 for short) metrics.
Dataset ogbn-arxiv ogbn-products ogbn-paper100M
Metrics Mi-F1 Ma-F1 Mi-F1 Ma-F1 Mi-F1 Ma-F1
w/o Ada 71.25 62.55 80.28 70.79 64.19 55.87
w/o Low 72.07 63.27 80.30 71.57 64.97 56.56
w/o LCS 71.49 62.70 80.42 71.04 64.45 55.98
w/o SC 73.02 64.01 81.58 72.23 65.46 57.19
ScaleGNN 73.25 64.20 81.82 72.47 65.69 57.42
The results demonstrate that each component is vital for improving ScaleGNN’s performance. w/o Ada variant experiences the largest performance drop across all datasets, with reductions of 1.77% in Micro-F1 and 1.53% in Macro-F1 on ogbn-arxiv, 1.16% and 1.54% on ogbn-products, and 1.34% and 1.42% on ogbn-paper100M, highlighting the importance of adaptive high-order feature fusion for capturing structural dependencies and mitigating information loss. w/o Low variant also shows notable degradation, especially on ogbn-arxiv and ogbn-paper100M, with Micro-F1 dropping by 0.95% and 0.56%, respectively, indicating that low-order enhanced feature aggregation is crucial for maintaining local information. w/o LCS variant sees moderate degradation, with Micro-F1 decreasing over three large-scale datasets, showing the effectiveness of high-order redundant feature masking in reducing redundancy and preventing excessive feature mixing. w/o SC, showing that it still falls short of the full-fledged ScaleGNN model despite being the best of all variants, demonstrating the important role of the LCS-based sparse constraints. It employs an attention-based relevance function, which allows ScaleGNN to adaptively assess and filter neighbors in a differentiable and task-driven manner, significantly improving its generalization and scalability in large-scale graphs.
Overall, ScaleGNN consistently outperforms all variants, confirming the significant contribution of each component. Adaptive high-order feature fusion is the most influential in improving performance, followed by low-order enhanced feature aggregation and high-order redundant feature masking. The combination of these mechanisms enables ScaleGNN to achieve superior results on various large-scale graph datasets.
V-F Trade-off Analysis: Performance vs. Efficiency (RQ5)
In this section, we evaluate the trade-off between efficiency and accuracy for ScaleGNN and ScaleGNNb by analyzing their runtime and Micro-F1 scores across diverse hop settings.
(a) ogbn-products
(b) ogbn-papers100M
Figure 4: The trade-off between efficiency and accuracy on large-scale graph datasets.
The runtime analysis reveals in Fig. 4 that ScaleGNN incurs a higher computational cost than ScaleGNNb across all hop settings. As the number of hops increases, both methods show rising runtimes, reflecting the increased complexity of message passing over larger neighborhoods. However, ScaleGNN consistently requires more computational resources, suggesting that its modeling choices introduce additional overhead.
ScaleGNN consistently outperforms ScaleGNNb in Micro-F1 scores. Both methods benefit from increasing hops, as it helps capture structural and semantic relationships. However, beyond a certain point, performance gains diminish due to redundant or noisy information from distant nodes. These results highlight the trade-off between accuracy and efficiency. While ScaleGNN achieves higher Micro-F1 scores, it comes with a significantly higher computational cost. In contrast, ScaleGNNb offers a more efficient alternative with a slight performance compromise. This suggests that model selection should be based on practical constraints: ScaleGNN is ideal for accuracy-focused tasks, while ScaleGNNb is better for resource-constrained environments. Balancing efficiency and accuracy remains crucial in large-scale graph learning.
(a) ogbn-products
(b) ogbn-papers100M
(c) ogbn-products
(d) ogbn-papers100M
Figure 5: The parameter sensitivity experimental results.
V-G Parameter Sensitivity Analysis (RQ6)
V-G1 The effect of hyperparameters $\beta$
We conduct a sensitivity analysis of $\beta$, which balances low-order and high-order feature aggregations, by varying its value from 0.1 to 0.9 and evaluating the Micro-F1 score (%). The results in Fig. 5 show that performance improves as $\beta$ increases, peaking at $\beta=0.5$. Beyond this point, further increases in $\beta$ lead to a gradual decline, likely due to overemphasizing high-order information, which may introduce noise or redundancy. These findings highlight the importance of $\beta$ to strike an optimal balance between local and global information aggregation.
V-G2 The effect of hyperparameters $\lambda_{1}$ and $\lambda_{2}$
We assess the sensitivity of $\lambda_{1}$ and $\lambda_{2}$ in the loss function by varying their values within {0.0005, 0.001, 0.005, 0.01, 0.05, 0.1} and evaluating Micro-F1. As shown in Fig. 5, ScaleGNN achieves the best performance at $\lambda_{1}=0.005$, with results improving up to this point and declining thereafter, reflecting the benefit of moderate regularization. In contrast, performance with respect to $\lambda_{2}$ is more stable, peaking at $\lambda_{2}=0.01$ with only slight variations, indicating its role in maintaining model stability. These results highlight the need for careful tuning: $\lambda_{1}$ requires precise adjustment for optimal results, while $\lambda_{2}$ ensures robustness. The loss terms weighted by $\lambda_{1}$ and $\lambda_{2}$ complement each other, balancing performance and efficiency.
VI Conclusion
We propose a novel scalable GNN model named ScaleGNN, tackling over-smoothing and scalability challenges in large-scale graphs. ScaleGNN adaptively aggregates informative high-order neighbors while suppressing redundancy through a Local Contribution Score (LCS)-based masking mechanism. Extensive experiments show that ScaleGNN outperforms state-of-the-art deep and scalable GNNs in both accuracy and efficiency. Future work includes extending it to heterogeneous graphs and optimizing large-scale training.